OpenStudy (anonymous):

Which is the most reasonable estimate? four and three-fourths times two and one-eighth minus seven and one-half

OpenStudy (jennychan12):

\[4\frac{ 3 }{ 4 }*2\frac{ 1 }{ 8 } - 7\frac{ 1 }{ 2 }\] 5*2 -7 10-7 3 actual: 2.59